Imagine that you are on a space walk outside a vessel coasting


through interstellar space. You have launched a reference probe that is


perfectly stationary relative to the ship, equipped with precision laser


velocity-measuring devices. A meteoroid approaches at 4.0 m/s, strikes


the ship, and bounces off in exactly the opposite direction at 2.0 m/s.


you grab the meteoroid, take it into the ship, put it into an inertial mass


meter, and find out that it masses 1.0 kg. The ship’s mass is 20,000 kg


without you on board. You know that the impact has caused the space


ship to start moving relative to the probe in the same direction the


meteoroid was originally travelling. What will the probe indicate as thmagnitude of the ship’s post-collision velocity vector?

1
Expert's answer
2021-11-30T17:44:28-0500

"mv_1=-mv_2+Mu,"

"u=\frac{m(v_1+v_2)}{M}=0.3~\frac{mm}s."